OpenStudy (shreenithaa):

13. Joining the frequent visitor program at the planetarium costs $5 per year, Write an equation that can be used to determine n, the number of visits per year for which the cost of being a member of the frequent visitor program is equal to the cost of not being a member. Tickets cost $11 for each student who is a member of the planetarium's frequent visitor program and $13 for each student who is not a member. 14 : Solve your equation from item 13. List a property of equality or provide an explanation for each step 15: Explain the meaning of the solution of the equation

OpenStudy (shreenithaa):

hi do u know the answer to the questions i need them by tomorrow morning at 5 thanks

OpenStudy (shreenithaa):

can u start out by answering the first then i can do it on my own and finish the rest

OpenStudy (jennyrlz):

hello :0

OpenStudy (shreenithaa):

hi

OpenStudy (shreenithaa):

my name is shree

OpenStudy (jennyrlz):

Pleasure to meet you ^.^

OpenStudy (shreenithaa):

u too

OpenStudy (shreenithaa):

do u get the question, i am stumped

OpenStudy (jennyrlz):

let me see... i will write as i read :)

OpenStudy (shreenithaa):

k thanks for helping

OpenStudy (jennyrlz):

planetarium program cost: 5 per yr n=visits per year perks for being a member: 2 dollars of ticket (non-memeber pirce: 13 dollars)

OpenStudy (jennyrlz):

ok i got this information from number thirteen

OpenStudy (shreenithaa):

k did u get the answer

OpenStudy (jennyrlz):

working it out xD

OpenStudy (shreenithaa):

sorry take ur time

OpenStudy (jennyrlz):

N = llx + 13y hmm im a little confused by the question but this is what i got :)

OpenStudy (jennyrlz):

X being the amount members visiting, and why the non members

OpenStudy (jennyrlz):

and y*

OpenStudy (shreenithaa):

that's what i got too but solving it was the hardest part i tried to search everywhere

OpenStudy (jennyrlz):

to solve for it we would need to make up an x and a y

OpenStudy (shreenithaa):

i think that x=13 and y=11 but how do you justify that that is your answer you know by using mathematical steps

OpenStudy (shreenithaa):

it says that they have to be equal thats why

OpenStudy (jennyrlz):

where does it say it has to be equal?

OpenStudy (shreenithaa):

it says it on number 13

OpenStudy (shreenithaa):

yup

OpenStudy (jennyrlz):

ohhhhhhhhhhhhhhhhhhhh i see, i just skimmed it XD

OpenStudy (shreenithaa):

its k

OpenStudy (jennyrlz):

let me re make the equation then :)

OpenStudy (shreenithaa):

k

OpenStudy (jennyrlz):

11n = 13n

OpenStudy (jennyrlz):

do you get it?

OpenStudy (shreenithaa):

yeah but when it says n is the number of total visitors, it just gets confusing

OpenStudy (jennyrlz):

ok let me explain it :)

OpenStudy (shreenithaa):

thnx

OpenStudy (jennyrlz):

n is the number of visitors, in other word total visitors not just members or vice versa. with me so far?

OpenStudy (shreenithaa):

guess so

OpenStudy (shreenithaa):

then how do you know how much n is

OpenStudy (jennyrlz):

one sec let me finish explaining that then we will get to that :)

OpenStudy (shreenithaa):

k

OpenStudy (jennyrlz):

hmm let me make an analogy (a fictional example to help relate)

OpenStudy (jennyrlz):

so if i were to try and get people to join a program i would try and show them the benefits, or how much money they can save correct?

OpenStudy (shreenithaa):

yup

OpenStudy (jennyrlz):

wait...

OpenStudy (jennyrlz):

i just confused my self... let me step back andre think this from the beggining

OpenStudy (shreenithaa):

sure

OpenStudy (jennyrlz):

ok.. after some quick thinking ive come up with a new equation

OpenStudy (jennyrlz):

5 + 11n = 13n

OpenStudy (jennyrlz):

wow it makes sense now :D

OpenStudy (shreenithaa):

yeah because of the +5 of the yearly thing but how about justifying it

OpenStudy (jennyrlz):

with this equation we get 2.5 visits per year which is wrong.. but we are on the right tracks now!

OpenStudy (jennyrlz):

lets make one more change to the equation

OpenStudy (jennyrlz):

remebered i said N was the total visits correct

OpenStudy (jennyrlz):

remember*

OpenStudy (shreenithaa):

s remembering

OpenStudy (jennyrlz):

this means that n can not be the same for both 11 and and 13

OpenStudy (shreenithaa):

yeah beacuse not the same amount of people have membership and not so ur right

OpenStudy (jennyrlz):

so the new equation is 5 +11x = N - 13y

OpenStudy (shreenithaa):

okay

OpenStudy (shreenithaa):

i just dont get how you will get n or x or y

OpenStudy (jennyrlz):

hmm same... but lets ee

OpenStudy (jennyrlz):

see

OpenStudy (shreenithaa):

it would be easier if we knew what n was wouldn't we

OpenStudy (jennyrlz):

hmm i just had an idea, let me try working it out on my side...

OpenStudy (shreenithaa):

sure

OpenStudy (shreenithaa):

i am sorry to hurry you but i need to go and could u type the answer i will check in the morning sorry i an stay for 10 minutes

OpenStudy (jennyrlz):

Alright ill see if this idea leads to the answer :D

OpenStudy (shreenithaa):

thns u have soooooooo much help

OpenStudy (shreenithaa):

what grade r u in

OpenStudy (jennyrlz):

12th

OpenStudy (shreenithaa):

whaa... i am in 8th

OpenStudy (jennyrlz):

but i finished highschool math last year xD

OpenStudy (jennyrlz):

im taking college lvl math now ^.^

OpenStudy (shreenithaa):

omg i have been waiting for colledge like forever

OpenStudy (shreenithaa):

i cant wait to go

OpenStudy (jennyrlz):

to think i passed calculus to be struggling on a middle school math problem u.u

OpenStudy (jennyrlz):

college is fun :d

OpenStudy (shreenithaa):

and hahaha well common core algebra 1

OpenStudy (shreenithaa):

do u come home from colledge or stay there

OpenStudy (jennyrlz):

go home

OpenStudy (jennyrlz):

xD

OpenStudy (shreenithaa):

oh guess that's why u r even here

OpenStudy (shreenithaa):

gotta go

OpenStudy (jennyrlz):

cya

OpenStudy (jennyrlz):

I GOT ITTTTTTT

OpenStudy (jennyrlz):

After many errors it all clicked!!!

OpenStudy (jennyrlz):

we where right the first time! well... errr third time or whatever it was

OpenStudy (jennyrlz):

the correct equation is 11n +5 = 13n

OpenStudy (jennyrlz):

what we need to prove is that buying membership and paing 11 dolars or just not being a member is the same in this equation

OpenStudy (jennyrlz):

so following the equation if you only go 2.5 times a year then it will cost you the same as if you buy membership and go 2.5 times a year

OpenStudy (jennyrlz):

11(2.5) = 27.5 =27.5 + 5 =32.5 13(2.5) = 32.5

OpenStudy (jennyrlz):

we did it!

OpenStudy (jennyrlz):

i was mislead by the impossibility of visiting 2.5 times, but that doesn't matter, the question is just asking for an equation that would make it theoretically possible for them them to result in the same price!!

OpenStudy (jennyrlz):

as for a property of equality...

OpenStudy (jennyrlz):

we used the subtraction property of equaltiy

OpenStudy (jennyrlz):

5+ 11n (-11n) = 13n (-11n)

OpenStudy (jennyrlz):

subtraction equality means that if we subtract the same number to both sides we still end up with the same equations for example the equation 5=5, if we subtract the same numbeer from both sides, it will still be true. 2-5 =5-2 = 3=3

OpenStudy (jennyrlz):

number 15 i think i answered by saying that you would have to visit 2.5 times etc....

OpenStudy (jennyrlz):

wow that was hard! That is the most thinking ive done in a while hahaha,